OpenStudy (anonymous):

Find all zeros write polynomial as product of linear equation determine x-intercepts X^4+25x^2+144 would like process not just answer thanks

hero (hero):

To find the zeroes of the expression, set the expression equal to zero, then solve for x: x^4+25x^2+144 = 0 x^4+9x^2+16x^2+144 = 0 x^2(x^2+9)+16(x^2+9) =0 (x^2+9)(x^2+16) = 0 x^2+9 = 0 x^2+16 = 0 The solution will be complex: x^2 = -9 x^2 = -16 x = +/-3i x = +/-4i

hero (hero):

Questions? Comments?

OpenStudy (anonymous):

Thankk youu